# On Artin’s conjecture on average and short character sums

### v1

@@ 1.1. Background and set-up
More precisely, if $a$ is an integer that is not $-1$ or a perfect square, let $b\in\mathbb{N}$ be the squarefree part of $a$ and let $h\geqslant 1$ be the largest integer such that $a$ is a perfect $h$-th power. We also define
@@
$N_{a}(x)=\#\{p\leqslant x\colon\leavevmode\nobreak\ a\textnormal{ is a primitive root modulo }p\},$
@@
  • if $b\not\equiv 1\pmod{4}$, then
@@
  • If $b\equiv 1\pmod{4}$, then
@@
#$\#\{|a|\leqslant y\colon\leavevmode\nobreak\ \limsup_{x\to\infty}N_{a}(x)<\infty\}\ll(\log y)^{2}.$
However, the methods of [3, 4] do not yield an asymptotic for $N_{a}(x)$ for almost all $a$ (and in fact the lower bounds are off from the conjectured magnitude by a factor of $(\log x)^{-1}$, since all the primes $p$ detected are such that $(p-1)/K$ has at most two prime factors for some small positive integer $K$).
@@ 1.3. Notation
We use $\sum^{\ast}_{r\pmod{q}}$ to denote summation over the primitive residue classes modulo $q$.
We denote by $\mathcal{X}_{q}$ the set of Dirichlet characters $\pmod{q}$ and by $\mathcal{X}_{q}^{*}$ the set of primitive Dirichlet characters $\pmod{q}$.
@@
Finally, we use #$\#{\mathcal{S}}$ to denote the cardinality of a finite set ${\mathcal{S}}$.
@@ 1.4. Acknowledgments
The authors thank Kaisa Matomäki for a helpful discussion on character sums.
@@ 2.1. Preliminaries
    $\frac{x}{2^{m}\log x}\ll\#\{n\leqslant x\colon\leavevmode\nobreak\ \Omega(n)=m\}\ll\frac{x}{2^{m}}\log\frac{x}{2^{m}}+1;$
@@
    #$\#\{n\leqslant x\colon\leavevmode\nobreak\ \Omega(n)=m\}\gg\frac{x}{2^{m}\log x}.$
@@
$\{n\leqslant x\colon\leavevmode\nobreak\ \Omega(n)=m\}=\{2^{m}\}.$
@@
For $103\log\log x\leqslant m\leqslant(\log x)/\log 2-1/$ 10, the bounds (i) and (ii) follow from a result of Nicolas [12] (see also [16, Part II, Equation (6.30) and Exercise 217], which in fact gives an asymptotic formula
#$\#\{n\leqslant x\colon\leavevmode\nobreak\ \Omega(n)=m\}=\left(C+O((\log y)^{-\eta})\right)y\log y$
where $y=x/2^{m}$, $\eta>0$ is some absolute constant, and the constant $C=0.3786\ldots$ is given by an explicit Euler products (note that $y\geqslant 2^{1/10}>1$).
@@
#$\#\{n\leqslant x\colon\leavevmode\nobreak\ \Omega(n)=m\}$$\geqslant\#\{2^{m}p\colon\leavevmode\nobreak\ p\leqslant x/2^{m}\}$
$=\pi(x/2^{m})\gg\frac{x}{2^{m}\log(x/2^{m})}\gg\frac{x}{2^{m}\log x},$
@@
${\mathcal{A}}_{w}(y)=\{a\leqslant y\colon\leavevmode\nobreak\ \Omega(a)=w\}.$
@@
since each choice of $b_{1},\ldots b_{k}$ corresponds to partitioning the multi-set of the $kw$ (not necessary distinct) primes dividing $b_{1}\cdots b_{k}$ into $k$ groups of $w$ primes. Using the elementary inequalities
@@
(2.3) $S_{k}(w,x,y)\ll wk^{kw+1}(x^{2}+y^{k})\left(\#{\mathcal{A}}_{w}(y)\right)^{k}.$
@@
(2.5) #$\begin{split}\#{\mathcal{E}}(w,x,y;\delta)&\ll\left(\#{\mathcal{A}}_{w}(y)\delta\right)^{-2k}y^{k+1}\left(\#{\mathcal{A}}_{w}(y)\right)^{k}wk^{kw+1}\\ &\ll kwy\left(\frac{k^{w}y}{\delta^{2}\#{\mathcal{A}}_{w}(y)}\right)^{k}.\end{split}$
@@
#$\#{\mathcal{A}}_{w}(y)\gg\frac{y}{2^{w}}\frac{1}{\log y}.$
@@
(2.7) #$\#{\mathcal{E}}(w,x,y;(\log y)^{-\lambda\log 2})\ll kwy\left(k^{w}(\log y)^{4\lambda\log 2}\right)^{k},$
@@
#$\#{\mathcal{E}}(w,x,y;(\log y)^{-\lambda\log 2})\ll x^{\xi},$
where
$\xi=2\frac{w\log\log x+4\lambda(\log 2)(\log\log y)}{\log y}+O\left(\frac{\log(k^{w}y)}{\log x}\right).$
Recalling our assumptions (1.5), (2.1), (2.6) and the estimate (2.8), we see that
$\frac{\log(k^{w}y)}{\log x}\ll\left(\log x\right)^{-0.48}.$
@@
(2.9) #$\#{\mathcal{E}}(w,x,y;(\log y)^{-\lambda\log 2})\leqslant x^{0.485}$
@@
$F_{\lambda}(y)=\#\{a\leqslant y\colon\leavevmode\nobreak\ \Omega(a)>\lambda\log\log y\}.$
@@
(2.10) $\begin{split}F_{\lambda}(y)&\leqslant\sum_{\lambda\log\log y<m\leqslant(\log y)/\log 2}\#\{a\leqslant y\colon\leavevmode\nobreak\ \Omega(a)=m\}\\ &\ll y\log y\sum_{m>\lambda\log\log y}2^{-m}+\log y\\ &\ll y(\log y)^{1-\lambda\log 2}+\log y\ll y(\log y)^{1-\lambda\log 2},\end{split}$
@@
#$\#{\mathcal{E}}(x,y)$ $\leqslant x^{0.485}\lambda\log\log y$
@@
Now let
${\mathcal{Q}}(x,y)=\{q\leqslant x\colon\leavevmode\nobreak\ q\not\in{\mathcal{E}}(x,y)\}.$
Then, for any natural numbers $w\leqslant\lambda\log\log y$ and $q\in{\mathcal{Q}}(x,y)$, we have
@@ 3.1. Moments of $N_{a}(x)$
Now, by Theorem 1.2, we see that there is a set $\mathcal{E}$ of size #$\#\mathcal{E}\ll x^{1/2}$ (say) such that for all primes $p\in[1,x]\setminus\mathcal{E}$ we have
@@
(3.4) $\frac{\varphi(p-1)}{p-1}\sum_{t\mid p-1}\frac{\mu(t)}{\varphi(t)}\sum_{\begin{subarray}{c}\chi\in{\mathcal{X}}_{p}\\ \mathrm{ord}(\chi)=t\end{subarray}}\chi(a)=\begin{cases}1&\text{if}\ a\in{\mathcal{G}}_{p},\\ 0,&\text{otherwise,}\\ \end{cases}$
@@
Separating the contribution of the principal character, corresponding to $t=1$, we arrive at
(3.5) $\sum_{p\in\mathcal{P}}\,\sum_{\begin{subarray}{c}-y\leqslant a\leqslant y\\ a\in{\mathcal{G}}_{p}\end{subarray}}1=\#([-y,y]\cap\mathbb{Z})\cdot\sum_{p\in{\mathcal{P}}}\frac{\varphi(p-1)}{p-1}+O(E)$
@@
$E=\sum_{p\in\mathcal{P}}\sum_{\begin{subarray}{c}t\mid p-1\\ t>1\end{subarray}}\frac{|\mu(t)|}{\varphi(t)}\sum_{\begin{subarray}{c}\chi\in{\mathcal{X}}_{p}^{*}\\ \mathrm{ord}(\chi)=t\end{subarray}}\left|\sum_{-y\leqslant a\leqslant y}\chi(a)\right|.$
@@
with $A$ given by (1.1), we can rewrite (3.5) as
@@
Since the group ${\mathcal{X}}_{p}$ is cyclic of order $p-1$, for each divisor $t\mid p-1$ there are $\varphi(t)$ characters $\chi\in{\mathcal{X}}_{p}^{*}$ with $\mathrm{ord}(\chi)=t$. Therefore we have
@@
${\mathcal{Q}}=\{(p_{1},p_{2})\in[1,x]^{2}\colon\leavevmode\nobreak\ (\log x)^{3D+2}\leqslant p_{1},p_{2}$ $\leqslant x,\leavevmode\nobreak\ p_{1}\neq p_{2},$
@@
#$\#\mathcal{Q}=\#\{(p_{1},p_{2})\in[1,x]^{2}\colon\leavevmode\nobreak\ (\log x)^{3D+2}\leqslant p_{1},p_{2}\leqslant x,\leavevmode\nobreak\ p_{1}$ $\neq p_{2}\}$
@@
The sums $S_{1}$ and $S_{2}$ are bounded symmetrically; let us bound $S_{1}$. Note that if $\chi_{1}$ is a non-principal character $\pmod{p_{1}}$ and $\chi_{0}$ is the principal character $\pmod{p_{2}}$, by the assumption $(p_{1},p_{2})\in\mathcal{Q}$ (which implies $p_{1}\not\in\mathcal{E}$ and $p_{2}\geq(\log x)^{3D+2}$), we have

### v2

@@ Abstract.
Key words and phrases:
primitive root, Artin’s conjecture on average, short character sums
2020 Mathematics Subject Classification:
11A07, 11L40, 11N25
@@ 1.1. Background and set-up
More precisely, if $a$ is an integer that is not 1-1 or a perfect square, let $b\in\mathbb{N}$ be the squarefree part of $a$ and let $h\geqslant 1$ be the largest integer such that $a$ is a perfect $h$-th power. We also define
@@
$N_{a}(x)=\#\{p\leqslant x\colon\penalty 10000\ a\textnormal{ is a primitive root modulo }p\},$
@@
  • if $b\not\equiv 1\hskip-3.41418pt\pmod{4}$, then
@@
  • If $b\equiv 1\hskip-3.41418pt\pmod{4}$, then
@@
$\{|a|\leqslant y\colon\penalty 10000\ \limsup_{x\to\infty}N_{a}(x)<\infty\}\ll(\log y)^{2}.$
However, the methods of [3, 4] do not yield an asymptotic for $N_{a}(x)$ for almost all $a$ (and in fact the lower bounds are off from the conjectured magnitude by a factor of $(\log x)^{-1}$, since all the primes $p$ detected are such that (p1)/K(p-1)/K has at most two prime factors for some small positive integer $K$).
@@ 1.3. Notation
We use $\sum^{\ast}_{r\hskip-3.41418pt\pmod{q}}$ to denote summation over the primitive residue classes modulo $q$.
We denote by $\mathcal{X}_{q}$ the set of Dirichlet characters $\hskip-3.41418pt\pmod{q}$ and by $\mathcal{X}_{q}^{*}$ the set of primitive Dirichlet characters $\hskip-3.41418pt\pmod{q}$.
@@
Finally, we use $\mathcal{S}$} to denote the cardinality of a finite set $\mathcal{S}$}.
@@ 1.4. Acknowledgments
The authors thank Kaisa Matomäki for a helpful discussion on character sums and the referee for the careful reading of the manuscript.
@@ 2.1. Preliminaries
    $\frac{x}{2^{m}\log x}\ll\#\{n\leqslant x\colon\penalty 10000\ \Omega(n)=m\}\ll\frac{x}{2^{m}}\log\frac{x}{2^{m}}+1;$
@@
    $\{n\leqslant x\colon\penalty 10000\ \Omega(n)=m\}\gg\frac{x}{2^{m}\log x}.$
@@
$\{n\leqslant x\colon\penalty 10000\ \Omega(n)=m\}=\{2^{m}\}.$
@@
For $3\log\log x\leqslant m\leqslant(\log x)/\log 2-1/10$, the bound (i), and thus (ii) in this range follows from a result of Nicolas [12] (see also [16, Part II, Equation (6.30) and Exercise 217], which in fact gives an asymptotic formula
(2.1) $\{n\leqslant x\colon\penalty 10000\ \Omega(n)=m\}=\left(C+O((\log y)^{-\eta})\right)y\log y$
where $y=x/2^{m}$, $\eta>0$ is some absolute constant, and the constant $C=0.3786\ldots$ is given by an explicit Euler product (note that $y\geqslant 2^{1/10}>1$). Indeed, we note that (2.1) always gives the desired upper bound (even in the regime when $y\ll 1$. To extract a lower bound from (2.1) we need to assume that $y\to\infty$. However since $2^{m}\in\{n\leqslant x\colon\penalty 10000\ \Omega(n)=m\}$ in the above range of $m$, for $y<\log\log x$ the lower bound is trivial, so we indeed get (i).
@@
$\#\{n\leqslant x\colon\penalty 10000\ \Omega(n)=m\}$ $\geqslant\#\{2^{m-1}p\colon\penalty 10000\ p\leqslant x/2^{m-1}\}$
$=\pi(x/2^{m-1})\gg\frac{x}{2^{m}\log(x/2^{m})}\gg\frac{x}{2^{m}\log x},$
@@
${\mathcal{A}}_{w}(y)=\{a\leqslant y\colon\penalty 10000\ \Omega(a)=w\}.$
@@
since each choice of $a_{1},\ldots,a_{k}$ corresponds to partitioning the multi-set of the $kw$ (not necessary distinct) primes dividing $b_{1}\cdots b_{k}$ into $k$ groups of $w$ primes. Using the elementary inequalities
@@
(2.4) $S_{k}(w,x,y)\ll wk^{kw+1}\left(x^{2}+y^{k}\right)\left(\#{\mathcal{A}}_{w}(y)\right)^{k}.$
@@
(2.6) $\begin{split}\#{\mathcal{E}}(w,x,y;\delta)&\ll\left(\#{\mathcal{A}}_{w}(y)\delta\right)^{-2k}y^{k+1}\left(\#{\mathcal{A}}_{w}(y)\right)^{k}wk^{kw+1}\\ &\ll kwy\left(\frac{k^{w}y}{\delta^{2}\#{\mathcal{A}}_{w}(y)}\right)^{k}.\end{split}$
@@
${\mathcal{A}}_{w}(y)\gg\frac{y}{2^{w}}\frac{1}{\log y}.$
@@
(2.8) ${\mathcal{E}}(w,x,y;(\log y)^{-\lambda\log 2})\ll kwy\left(k^{w}(\log y)^{4\lambda\log 2}\right)^{k},$
@@
${\mathcal{E}}(w,x,y;(\log y)^{-\lambda\log 2})\ll x^{\xi},$
where, using the upper bound on $k$ from (2.9), we conclude
$\xi=2\frac{w\log\log x+4\lambda(\log 2)(\log\log y)}{\log y}+O\left(\frac{\log(kwy)}{\log x}\right).$
Recalling our assumptions (1.5), (2.2), (2.7) and the estimate (2.9) again, we see that
$\frac{\log(kwy)}{\log x}\ll\left(\log x\right)^{-0.48}.$
@@
(2.10) ${\mathcal{E}}(w,x,y;(\log y)^{-\lambda\log 2})\leqslant x^{0.485}$
@@
$F_{\lambda}(y)=\#\{a\leqslant y\colon\penalty 10000\ \Omega(a)>\lambda\log\log y\}.$
@@
(2.11) $\begin{split}F_{\lambda}(y)&\leqslant\sum_{\lambda\log\log y<m\leqslant(\log y)/\log 2}\#\{a\leqslant y\colon\penalty 10000\ \Omega(a)=m\}\\ &\ll y\log y\sum_{m>\lambda\log\log y}2^{-m}+\log y\\ &\ll y(\log y)^{1-\lambda\log 2}+\log y\ll y(\log y)^{1-\lambda\log 2},\end{split}$
@@
$\#{\mathcal{E}}(x,y)$ $\leqslant x^{0.485}\lambda\log\log y$
@@
Then, for any natural numbers $w\leqslant\lambda\log\log y$ and $q\leqslant x$ with $q\not\in{\mathcal{E}}(x,y)$, we have
@@ 3.1. Moments of $N_{a}(x)$
Now, by Theorem 1.2, we see that there is a set $\mathcal{E}$ of size $\mathcal{E}\ll x^{1/2}$ (say) such that for all primes $p\in[1,x]\setminus\mathcal{E}$ we have
@@
[…]
@@
Separating the contribution of the principal character, corresponding to $t=1$ in (3.4), we arrive at
(3.5) $\sum_{p\in\mathcal{P}}\,\sum_{\begin{subarray}{c}-y\leqslant a\leqslant y\\ a\in{\mathcal{G}}_{p}\end{subarray}}1=\#([-y,y]\cap\mathbb{Z})\cdot\sum_{p\in{\mathcal{P}}}\frac{\varphi(p-1)}{p-1}+O(E+y\log y)$
@@
$E=\sum_{p\in\mathcal{P}}\sum_{\begin{subarray}{c}t\mid p-1\\ t>1\end{subarray}}\frac{|\mu(t)|}{\varphi(t)}\sum_{\begin{subarray}{c}\chi\in{\mathcal{X}}_{p}^{*}\\ \mathrm{ord}(\chi)=t\end{subarray}}\left|\sum_{-y\leqslant a\leqslant y}\chi(a)\right|,$
where the $O(y\log y)$ term comes from the contribution
$\sum_{p\leqslant y}\left(2y/p+O(1)\right)=2y\sum_{p\leqslant y}1/p+O(1)\ll y\log y$
of the principal characters modulo $p$ on values $a\equiv 0\hskip-3.41418pt\pmod{p}$.
@@
with AA given by (1.1), we can rewrite (3.5) as
@@
Since the group ${\mathcal{X}}_{p}$ is cyclic of order p1p-1, for each divisor $t\mid p-1$ there are $\varphi(t)$ characters $\chi\in{\mathcal{X}}_{p}^{*}$ with $\mathrm{ord}(\chi)=t$. Therefore we have
@@
${\mathcal{Q}}=\{(p_{1},p_{2})\in[1,x]^{2}\colon\penalty 10000\ (\log x)^{3D+2}\leqslant p_{1},p_{2}$ $\leqslant x,\penalty 10000\ p_{1}\neq p_{2},$
@@
$\#\mathcal{Q}=\#\{(p_{1},p_{2})\in[1,x]^{2}\colon\penalty 10000\ (\log x)^{3D+2}\leqslant p_{1},p_{2}\leqslant x,\penalty 10000\ p_{1}$ $\neq p_{2}\}$
@@
The sums $S_{1}$ and $S_{2}$ are bounded symmetrically; let us bound $S_{1}$. Note that if $\chi_{1}$ is a non-principal character $\hskip-3.41418pt\pmod{p_{1}}$ and $\chi_{0}$ is the principal character $\hskip-3.41418pt\pmod{p_{2}}$, by the assumption $(p_{1},p_{2})\in\mathcal{Q}$ (which implies $p_{1}\not\in\mathcal{E}$ and $p_{2}\geq(\log x)^{3D+2}$), we have